OpenStudy (anonymous):

Add. Simplify by collecting like radical terms if possible. 6sqrt8 + 5sqrt50=

OpenStudy (anonymous):

\[6\sqrt 8 + 5\sqrt50=\]

OpenStudy (anonymous):

37sqrt2

OpenStudy (anonymous):

Need an explanation on how to do this?

OpenStudy (anonymous):

you can "simplify" via \[\sqrt{8}=2\sqrt{2}\] \[\sqrt{50}=5\sqrt{2}\] so you have \[12\sqrt{2}+25\sqrt{2}=37\sqrt{2}\]

OpenStudy (anonymous):

I got it now! Thanks guys! =)

OpenStudy (anonymous):

Ok, you're very welcome.. satellite laid it out pretty well :)